OpenStudy (anonymous):

Factor the polynomial completely and find al zeros P(x)=x^6-729 Please help!

OpenStudy (anonymous):

\[P(x)= x^{6}-729\]

OpenStudy (welshfella):

x^6 - 729 = (x^3)^2 - 27^2 = (x^3 - 27)(x^3 + 27)

OpenStudy (welshfella):

now you can factor each parentheses using the cube identities a^3 + b^3 = (a + b)(a^2 - ab + b^2) and a^3 - b^3 = (a - b)(a^2 + ab + b^2)

OpenStudy (welshfella):

putting a^3 = x^3 and b^3 = 27 where a= x and b = 3

OpenStudy (welshfella):

to find the zeros let (x^3 - 27)( x^3 + 27) = 0 and solve for x

OpenStudy (anonymous):

\[(x+3)\times (x^{2}-3x+9)\times (x-3)\times (x ^{2}+3x+9)\]

OpenStudy (anonymous):

is that right?.. for the zeros.. should i just use the the quadratic equations to get the remained zeros?

OpenStudy (welshfella):

yes that's it

OpenStudy (anonymous):

AAAHH thank you!

OpenStudy (welshfella):

hmm maybe its better to find the zeroes using the last expression x = 3 and -3 are obvious roots to find the other 4 roots you need to solve x^2 - 3x + 9 = 0 and x^2 + 3x + 9 = 0

OpenStudy (anonymous):

could i solve them with the quadratic equation?

OpenStudy (welshfella):

they will all be complex roots

OpenStudy (welshfella):

quadratic formula yes

OpenStudy (anonymous):

cool! thank you.. you were really helpful!

OpenStudy (welshfella):

yw